OpenStudy (anonymous):

A playground is being designed where children can interact with their friends in certain combinations. If there is 1 child, there can be 0 interactions. If there are 2 children, there can be 1 interaction. If there are 3 children, there can be 4 interactions. If there are 4 children, there can be 9 interactions. How many interactions will there be for 6 children? A) 11 B) 16 C) 25 D) 36

OpenStudy (anonymous):

Im pretty sure it's 25

OpenStudy (princeharryyy):

how did u do that @savannawhiddenUSMC

OpenStudy (shinalcantara):

yes it's 25

OpenStudy (princeharryyy):

when u go through reality this isn't possible although the function is such that the ans gives 25 @shinalcantara nd @savannawhiddenUSMC

OpenStudy (princeharryyy):

In real the total handshakes would be 6*5/2 = 15 handshakes

OpenStudy (princeharryyy):

for this absurd ques ... 1 people 0 hskes.. 2 people 1*1 =1 3 people 2*2 =4 4people 3*3 =9 5 people 4*4 =16 6 people 5*5 =25

OpenStudy (shinalcantara):

trying to figure out the pattern of the problem, we will get the interaction to be (x-1)^2 where x is the number of children

OpenStudy (anonymous):

So Im right? ..

OpenStudy (shinalcantara):

yep

OpenStudy (princeharryyy):

yep u are don't just do it in real life u'll fail in counting @savannawhiddenUSMC

OpenStudy (princeharryyy):

for this question it was ohk

OpenStudy (anonymous):

thanks :) @princeharryyy

OpenStudy (princeharryyy):

u r welcome @savannawhiddenUSMC